# Social welfare optimisation in well-mixed and structured populations

### v1

@@ Abstract
Research on promoting cooperation among autonomous, self–regarding agents has often focused on the bi–objective optimization problem: minimizing the total incentive cost while maximising the frequency of cooperation. However, the optimal value of social welfare under such constraints remains largely unexplored. In this work, we hypothesise that achieving maximal social welfare is not guaranteed at the minimal incentive cost required to drive agents to a desired cooperative state. To address this gap, we adopt to a single–objective approach focused on maximising social welfare, building upon foundational evolutionary game theory models that examined cost efficiency in finite populations, in both well-mixed and structured population settings. Our analytical model and agent–based simulations who how different interference strategies, including rewarding local versus global behavioural patterns, affect social welfare an dynamics of cooperation. Our results reveal a significant gap in the per–individual incentive cost between optimising for pure cost efficiency or cooperation frequency and optimising for maximal social welfare. Overall, our findings indicate that incentive design, policy, and benchmarking in multi‑agent systems and human societies should prioritise welfare‑centric objectives over proxy targets of cost or cooperation frequency.
@@ 1 Introduction
The evolution of cooperation has long been a central puzzle in evolutionary biology, social sciences and multi–agent systems [sigmund2010calculus, perc2017statistical]. While classical evolutionary theory emphasizes the survival of the fittest–which in many strategic settings (e.g., the Prisoner’s Dilemma) often corresponds to more selfish behaviour–cooperative behaviour is nevertheless pervasive among both humans and animals. This apparent contradiction has motivated extensive research into the mechanisms that enable and explain how cooperation to emerge and persist in social dilemmas [nowak2006five].
To address this puzzle of cooperation [perc2017statistical, rand2011evolution], numerous mechanisms have been proposed to account for the emergence of cooperation, including kin and group selection [hamilton1964genetical, traulsen2006evolution], direct and indirect reciprocity [nowak:2005:nature, axelrod1981evolution], reward and punishment (incentives) [fehr2000cooperation, chen2015first] and structured populations [szabo2007evolutionary]. A particularly important mechanism examined in recent years involves an external institution that seeks to steer the population towards greater cooperation by selectively investing in individuals [chen2015first, duong2023cost]. For example, such an institution may provide rewards to cooperators, either at the global population level or in a more localized, neighbourhood–based manner. However, existing research often overlooks social welfare and fails to address the delicate balance between fostering cooperation and optimizing welfare.
In well–mixed populations, Han and Tran–Thanh (2018) [han2018interference] showed that a decision–maker can conditionally reward cooperators based on population composition to guarantee a desired cooperation level while minimizing interference cost. Their framework formulates a bi–objective optimisation problem: maximize cooperation frequency while minimizing institutional investment. Later, Duong and Han (2021) [duonghan2021cost] provided a rigorous stochastic analysis of institutional incentives, characterizing optimal reward/punishment schemes under different selection intensities and identifying sharp phase transition phenomena in cost efficiency.
@@
Despite these advances, prior works mainly focus on either (i) maximising cooperation prevalence or (ii) minimising the institutional cost [duonghan2021cost, wang2019exploring]. What is missing is a direct optimisation of social welfare, defined as:
@@
From a societal or institutional perspective, social welfare is ultimately the most meaningful objective: cooperation is valuable insofar as it generates net benefits relative to the cost of enforcing it [karsu2015inequity, nyman2006efficiency, HanInterface2024Welfare, HAN2026_social_welfare]. Building on existing work, this paper introduces social welfare maximization into the study of cost–efficient external interference. We address the following research questions:
  • RQ1: How does optimizing social welfare change the optimal interference strategies in well–mixed populations? Does welfare maximization require less, more, or differently patterned investment compared to minimising cost alone?
@@
Our contributions are summarized as follows:
@@
Our overall approach can be summarised in Figure 1. By introducing social welfare into evolutionary models of institutional incentives, this work offers a more realistic and policy–relevant framework for understanding how cooperation can be engineered in biological, social, and artificial systems. Our unified perspective bridges analytical well–mixed models and simulation–based structured models, shedding new light on the design of welfare–maximizing intervention schemes.
@@ 2 Literature review
A prominent line of work examines external institutional influence, where a central authority invests resources to steer population behaviour [duong2023cost, sigmund2001reward, cimpeanu2021cost]. Early models typically assumed well–mixed populations and studied how rewards, punishments, or other incentive schemes affect the stability of cooperation. Han and Tran–Thanh (2018) introduced a conditional investment mechanism in which an external controller (EC) allocates rewards depending on the current population state [han2018interference]. Their analysis showed that the EC can guarantee a desired cooperation level while minimizing the total interference cost, formally framing the problem as a bi–objective optimisation between maximizing cooperation and minimizing investment. Building on this, Duong and Han (2021) provided a more comprehensive stochastic treatment, deriving closed–form results for cost efficiency under various selection intensities and revealing nontrivial threshold behaviours in the optimal interference levels [duonghan2021cost].
Beyond well–mixed populations, spatial and network structures introduce additional complexity [santos2006cooperation, perc2013evolutionary, nowak1992evolutionary]. Real systems often involve locality constraints—agents interact more frequently with a subset of neighbors rather than the entire population. Han et al. (2018) extended institutional incentive models to structured populations using agent–based simulations on grids [han2018fostering]. They demonstrated that local interference strategies, which condition investment on neighborhood–level information, can outperform global strategies in terms of both cooperation levels and cost efficiency. These results highlight the importance of spatial heterogeneity in designing external incentive mechanisms.
@@
In summary, while prior works have laid a solid foundation for understanding institutional incentives in evolutionary dynamics, a unified welfare–based analysis across both well–mixed and structured populations is still missing. Addressing this gap is essential not only for theoretical completeness but also for practical relevance, especially in contexts such as public policy, distributed AI systems, and resource allocation, where maximizing net societal benefit is a central goal.
@@ 3 Models and Methods
This section reviews the original Evolutional Game Theory model for previous optimization problems in [duonghan2021cost], and proposes a new methodology to compute and assess the overall social welfare of a well–mixed population. Furthermore, it also recaps the agent–based simulation setup extending from [han2018fostering], where it briefly goes over the simulation process, the addition of social welfare as a metric to analyse, and the different intervention strategies that this work covers.
@@ 3.1 Problem setup
We consider a well–mixed, finite population of $N$ players, who interact with each other using cooperation dilemmas, such as the Donation Game or Public Goods Game. Let $\Pi_{C}(i)\text{ and }\Pi_{D}(i)$ be the expected payoffs of a cooperative player (C–player) and a defective player (D–player) in state $S_{i}$ of the population, respectively. Note that $S_{i}$ represents a state where the population has $i$ C–players. We assume $\Pi_{C}(i)\text{ and }\Pi_{D}(i)$ monotonically depend on $i$, and the difference $\delta=\Pi_{C}(i)-\Pi_{D}(i)$ is a negative metric, independent of $i$.
@@
To reward a cooperator (respectively, punish a defector), the institution has to spend an amount $\theta/a$ (respectively, $\theta/b$), such that the payoff of the targeted individual increases (or decreases) by $\theta$, where $a,b>0$ denote the efficiency ratios of reward and punishment, respectively.
@@
$\min\Big(\frac{i}{a},\frac{N-i}{b}\Big)\theta.$
The central question is thus: what is the minimal incentive intensity $\theta$ that ensures a prescribed long–run level of cooperation while minimising the total institutional cost?
@@
$\theta_{j}=\min\Big(\frac{j}{a},\frac{N-j}{b}\Big)\theta.$
@@
$E_{\text{mix}}(\theta)=\frac{\theta}{2}\sum_{j=1}^{N-1}(n_{1j}+n_{N-1,j})\min\Big(\frac{j}{a},\frac{N-j}{b}\Big).$ (2)
@@
$E_{r}(\theta)=\frac{\theta}{2}\sum_{j=1}^{N-1}(n_{1j}+n_{N-1,j})\frac{j}{a},\qquad E_{p}(\theta)=\frac{\theta}{2}\sum_{j=1}^{N-1}(n_{1j}+n_{N-1,j})\frac{N-j}{b}.$ (3)
We note that, unlike the separate reward and punishment schemes, the efficiency ratios $a$ and $b$ directly affect the structure of the mixed incentive cost $E_{\text{mix}}(\theta)$ through their interaction in the minimum operator, making the optimisation problem substantially more involved.
@@ 3.2 Analytical calculation
We will derive an expression for the expected total social welfare across population states, $SW(\theta)(\theta)$, under the influence of a uniform institutional reward $\theta$ for cooperative players. The social welfare in state $S_{i}$ (where $i$ is the number of cooperators) is the total payoff received by all players ($P_{i}$) minus the total institutional incentives paid ($\theta_{i}$). We assume the reward is funded from an external budget, meaning the institution’s spending does not reduce the population’s baseline payoff.
@@
To evaluate the expected population-level social welfare ($SW(\theta)(\theta)$), we aggregate $SW_{i}(\theta)(\theta)$ over all transient states $S_{i}$ ($i=1$ to N1N-1), weighted by the expected visitation rates ($\frac{n_{1,i}+n_{N-1,i}}{2}$) derived from the Fundamental Matrix ($\mathcal{N}$) defined in [duonghan2021cost].
@@ 3.3.1 Prisoner’s Dilemma on Square Lattice Networks
In mathematical analysis, we used well–mixed network to have a general observation. However, the connection between individuals in population in reality often different from that in literature research, we apply squared lattice – a widely used network structure in many approaches for evolutionary games (see [han2018fostering]) – to create the population structure of agents. We also adopted the agent–based model similar to the recent setup of experiments on cooperation in lab for easy comparation.
Specifically, we use network with size $Z=L\times L$. At the beginning, each agent is assigned either as a cooperator (C) or defector (D) with the same probability. We use the one–shot Prisoner’s Dilemma game to model the interaction between agents, where R is the reward for the cooperation (penalty P) and the denial in cooperation gives the defector the temptation T and the fool payoff S. We adopt the scaled payoff matrix of the PD [han2018fostering]: $T=b,R=1,P=S=0$ (with $1<b\leq 2$).
At each time step or generation, we allow each agent play the PD the same as [han2018fostering] in which it interacts with its four neighbours and decides whether to change the strategy or not based on the neighbour’s highest scores. We follow the deterministic, standard evolutionary process in order to capture the cost of different interference strategies for the social welfare efficiency analysis. In addition, we also conduct experiments on a stochastic update rule to prove the correctness of hypothesis. Instead of choosing the highest scores among neighbors, agent A with score $f_{A}$ has a Fermi probability that it copies the strategy from agent B (randomly chosen among neighbours) with score $f_{B}$ with the amplitude of noise $K$  [traulsen2006stochastic, szabo2007evolutionary]: $(1+e^{(f_{A}-f_{B})/K})^{-1}$.
We run the evolutionary process until it reaches either a steady state or a repeating cycle. To ensure a consistent and fair comparison, each simulation is executed for 50 generations. To further enhance accuracy, the final results for each parameter setting are computed by averaging 10 independent runs. We do not consider mutations or explorations in this work similar to [han2018fostering].
@@ 3.3.2 Social Welfare Efficiency
As stated before, we argue that minimising the cost of interference while fostering the level of cooperation in a structured network can not maximising the social welfare. We aim to study how to achieve the efficiency in social welfare and also reach the optimisation of interference cost.
Investors pay a cost $\theta>0$ for a cooperator (to the external decision–maker/investor). In particular, we investigate whether global interference strategies (where investments are triggered based on network level information) or their local counterparts (where investments are based on local neighbourhood information) lead to successful behaviour with better social welfare efficiency. To do so, we conduct two main classes of interference strategies the same as [han2018fostering] to capture the social welfare efficiency.
3.3.3 Global versus Local Intervention
Simplifying from [han2018fostering], we consider two external intervention schemes: global (population–composition–based – POP) and local (neighborhood–based – NEB).
@@ 4 Results
The first two parts of this section summarizes the results of the mathematical analysis of the objective function $SW$ w.r.t both $\theta$ and $\beta$, including its variability and limit behaviours, as well as an approximation algorithm to calculate the optimal solution and simulations for various cases of the Donation Game. The latter part of this section discusses the results of the agent–based simulation on a square lattice network, where it goes over the configuration of the experiments for both global and local intervention schemes, and utilises different figures as different angles to interpret the result.
@@ 4.1.3 Algorithm to Approximate the Optimal Incentive for Maximum Social Welfare
Below is the pseudocode of our simple local search algorithm to find an approximation of the optimal incentive $\theta^{*}$ maximizing social welfare.
@@
4.1.4 Influence of Selection Intensity ($\beta$)
  • Neutral Selection Limit ($\beta\rightarrow 0^{+}$)
  • Strong Selection Limit ($\beta\rightarrow+\infty$)
    $\lim_{\beta\rightarrow+\infty}\frac{f(x)}{g(x)}$ $=H_{N}+1$

### v2

@@ Abstract
Research on promoting cooperation among autonomous, self–regarding agents has often focused on the bi–objective optimization problem: minimizing the total incentive cost while maximising the frequency of cooperation. However, the optimal value of social welfare under such constraints remains largely unexplored. In this work, we hypothesise that achieving maximal social welfare is not guaranteed by the minimal incentive cost required to drive agents to a desired cooperative state. To address this gap, we adopt to a single–objective approach focused on maximising social welfare, building upon foundational evolutionary game theory models that examined cost efficiency in finite populations, in both well-mixed and structured population settings. Our analytical model and agent–based simulations show how different interference strategies, including rewarding local versus global behavioural patterns, affect social welfare and dynamics of cooperation. Our results reveal a significant gap in the per–individual incentive cost between optimising for pure cost efficiency or cooperation frequency and optimising for maximal social welfare. Overall, our findings indicate that incentive design, policy, and benchmarking in multi‑agent systems and human societies should prioritise welfare‑centric objectives over proxy targets of cost or cooperation frequency.
@@ 1 Introduction
The evolution of cooperation has long been a central puzzle in evolutionary biology, social sciences and multi–agent systems [sigmund2010calculus, perc2017statistical, han2022emergent, paiva2018engineering, tuyls2007evolutionary]. While classical evolutionary theory emphasizes the survival of the fittest— which in many strategic settings (e.g., the Prisoner’s Dilemma) often corresponds to more selfish behaviour—cooperative behaviour is nevertheless pervasive among both humans and animals. This apparent contradiction has motivated extensive research into the mechanisms that enable and explain how cooperation emerges and persists in social dilemmas [nowak2006five].
To address this puzzle of cooperation [perc2017statistical, rand2011evolution], numerous mechanisms have been proposed to account for the emergence of cooperation, including kin and group selection [hamilton1964genetical, traulsen2006evolution], direct and indirect reciprocity [nowak:2005:nature, axelrod1981evolution], reward and punishment (incentives) [fehr2000cooperation, chen2015first] and structured populations [szabo2007evolutionary]. A particularly important mechanism examined in recent years involves an external institution that seeks to steer the population towards greater cooperation by selectively investing in individuals [chen2015first, duong2023cost, wang2022incentive, sigmund2001reward]. For example, such an institution may provide rewards to cooperators, either at the global population level or in a more localized, neighbourhood–based manner. However, existing research often overlooks social welfare and fails to address the delicate balance between fostering cooperation and optimizing social welfare.
In well–mixed populations, Han and Tran–Thanh (2018) [han2018interference] showed that a decision–maker can conditionally reward cooperators based on population composition to guarantee a desired cooperation level, while minimizing interference cost. Their framework formulates a bi–objective optimisation problem: maximize cooperation frequency while minimizing institutional investment. Later, Duong and Han (2021) [duonghan2021cost] provided a rigorous stochastic analysis of institutional incentives, characterizing optimal reward/punishment schemes under different selection intensities and identifying sharp phase transition phenomena in cost efficiency.
@@
Despite these advances, prior works mainly focus on either (i) maximising cooperation prevalence or (ii) minimizing the institutional cost [duonghan2021cost, wang2019exploring]. What is missing is a direct optimisation of social welfare, defined as:
@@
From a societal or institutional perspective, social welfare is ultimately the most meaningful objective: cooperation is valuable insofar as it generates net benefits relative to the cost of enforcing it [karsu2015inequity, nyman2006efficiency, HanInterface2024Welfare, HAN2026_social_welfare]. This paper introduces social welfare maximization into the study of cost–efficient external interference. We address the following research questions:
  • RQ1: How does optimizing social welfare change the optimal interference strategies in well–mixed populations? Does welfare maximization require less, more, or differently patterned investment compared to minimizing cost alone?
@@
Our key contributions are summarized as follows:
@@
Our overall approach can be summarized in Figure 1. By introducing social welfare into evolutionary models of institutional incentives, this work offers a more realistic and policy–relevant framework for understanding how cooperation can be engineered in biological, social, and artificial systems. Our unified perspective bridges analytical well–mixed models and simulation–based structured models, shedding new light on the design of welfare–maximizing intervention schemes.
@@ 2 Literature review
A prominent line of work examines external institutional influence, where a central authority invests resources to steer population behaviour [duong2023cost, sigmund2001reward, cimpeanu2021cost]. Early models typically assumed well–mixed populations and studied how rewards, punishments, or other incentive schemes affect the stability of cooperation. Han and Tran–Thanh (2018) introduced a conditional investment mechanism in which an external controller (EC) allocates rewards depending on the current population state [han2018interference]. Their analysis showed that the EC can guarantee a desired cooperation level while minimizing the total interference cost, formally framing the problem as a bi–objective optimisation between maximizing cooperation and minimizing investment. Building on this, Duong and Han (2021) provided a more comprehensive stochastic treatment, deriving closed–form results for cost efficiency under various selection intensities and revealing non-trivial threshold behaviours in the optimal interference levels [duonghan2021cost].
Beyond well–mixed populations, spatial and network structures introduce additional complexity [santos2006cooperation, perc2013evolutionary, nowak1992evolutionary]. Real systems often involve locality constraints—agents interact more frequently with a subset of neighbours rather than the entire population. Han et al. (2018) extended institutional incentive models to structured populations using agent–based simulations on grids [han2018fostering]. They demonstrated that local interference strategies, which condition investment on neighbourhood–level information, can outperform global strategies in terms of both cooperation levels and cost efficiency. These results highlight the importance of spatial heterogeneity in designing external incentive mechanisms.
@@
In summary, while prior works have laid a solid foundation for understanding institutional incentives in evolutionary dynamics, a unified social welfare–based analysis across both well–mixed and structured populations is still missing. Addressing this gap is essential not only for theoretical completeness but also for practical relevance, especially in contexts such as public policy, distributed AI systems, and resource allocation, where maximizing net societal benefit is a central goal.
@@ 3 Models and Methods
This section reviews the original Evolutional Game Theory model for previous optimization problems in [duonghan2021cost], and derives the social welfare function in a well–mixed population setting. Furthermore, it describes the agent–based simulation setup extending from [han2018fostering], including the simulation process, the formulation of social welfare, and the different intervention strategies that this work will study.
@@ 3.1 Problem setup
We consider a well–mixed, finite population of $N$ players, who interact with each other using cooperation dilemmas, such as the Donation Game or Public Goods Game. Let $\Pi_{C}(i)\text{ and }\Pi_{D}(i)$ be the expected payoffs of a cooperative player (C–player) and a defective player (D–player) in state $S_{i}$ of the population, respectively. $S_{i}$ represents a state where the population has $i$ C–players.
@@
To reward a cooperator (respectively, punish a defector), the institution has to spend an amount $\theta/a$ (respectively, $\theta/\hat{a}$), such that the payoff of the targeted individual increases (or decreases) by $\theta$, where $a,\hat{a}>0$ denote the efficiency ratios of reward and punishment, respectively.
@@
$\min\Big(\frac{i}{a},\frac{N-i}{\hat{a}}\Big)\theta.$
The central question is thus: what is the minimal incentive intensity $\theta$ that ensures a prescribed long–run level of cooperation while minimizing the total institutional cost?
@@
$\theta_{j}=\min\Big(\frac{j}{a},\frac{N-j}{\hat{a}}\Big)\theta.$
@@
[…]
@@
$E_{r}(\theta)=\frac{\theta}{2}\sum_{j=1}^{N-1}(n_{1j}+n_{N-1,j})\frac{j}{a},\qquad E_{p}(\theta)=\frac{\theta}{2}\sum_{j=1}^{N-1}(n_{1j}+n_{N-1,j})\frac{N-j}{\hat{a}}$. (3)
We note that, unlike the separate reward and punishment schemes, the efficiency ratios $a$ and $b$ directly affect the structure of the mixed incentive cost $E_{\text{mix}}(\theta)$ through their interaction in the minimum operator, making the optimisation problem more complex.
@@ 3.2 Analytical calculation
We first derive an expression for the expected total social welfare across population states, $SW(\theta)$, under the influence of a institutional reward $\theta$ for cooperative players. The social welfare in state $S_{i}$ (where $i$ is the number of cooperators) is the total payoff received by all players ($P_{i}$) minus the total institutional cost paid ($\theta_{i}$).
@@
To evaluate the expected population-level social welfare, we aggregate $SW_{i}(\theta)$ over all transient states $S_{i}$ ($i=1$ to N1N-1), weighted by the expected visitation rates ($\frac{n_{1,i}+n_{N-1,i}}{2}$) derived from the Fundamental Matrix ($\mathcal{N}$) defined in [duonghan2021cost].
@@ 3.3.1 Prisoner’s Dilemma on Square Lattice Networks
Beyond the well-mixed population assumption above, we distribute agents on a square lattice, a network structure widely used in evolutionary game studies [szabo2007evolutionary, han2018fostering]. Specifically, we use a network with size $Z=L\times L$. At the beginning, each agent is assigned either as a cooperator (C) or defector (D) with the same probability. We use the one–shot Prisoner’s Dilemma game to model the interaction between agents, where R is the reward for the cooperation (penalty P) and the denial in cooperation gives the defector the temptation T and the fool payoff S. We adopt the scaled payoff matrix of the PD [han2018fostering]: $T=b,R=1,P=S=0$ (with $1<b\leq 2$).
At each time step or generation, each agent plays the PD with its four neighbours and decides whether to change the strategy or not based on the neighbour’s highest scores. We follow the deterministic, standard evolutionary process in order to capture the cost of different interference strategies for the social welfare efficiency analysis. In addition, we also conduct experiments on a stochastic update rule to prove the correctness of hypothesis. Instead of choosing the highest scores among neighbours, agent A with score $f_{A}$ has a Fermi probability that it copies the strategy from agent B (randomly chosen among neighbours) with score $f_{B}$ with the amplitude of noise $K$  [traulsen2006stochastic, szabo2007evolutionary]: $(1+e^{(f_{A}-f_{B})/K})^{-1}$.
We run the evolutionary process until it reaches either a steady state or a repeating cycle. To ensure a consistent and fair comparison, each simulation is executed for 50 generations. To further enhance accuracy, the final results for each parameter setting are computed by averaging 10 independent runs. We do not consider mutations or explorations in this work, for a convenient comparison with [han2018fostering].
@@ 3.3.2 Social Welfare Efficiency
As noted above, we hypothesise that minimising interference costs while increasing cooperation does not necessarily maximise population social welfare. We demonstrate this in the present study.
Investors pay a cost $\theta>0$ for a cooperator (to the external decision–maker/investor). In particular, we investigate whether global interference strategies (where investments are triggered based on population level information) or their local counterparts (where investments are based on local neighbourhood information) lead to successful behaviour with better social welfare efficiency. To do so, we analyse two main classes of interference strategies [han2018fostering] to for their capacity to optimise social welfare: i) global (population–composition–based – POP) and ii) local (neighborhood–based – NEB).
@@ 4 Results
The first two parts of this section summarise the results of the mathematical analysis of the objective function $SW$ with respect to both $\theta$ and $\beta$, including its variability and limiting behaviours, as well as an approximation algorithm to compute the optimal solution and simulations for various cases of the Donation Game. The latter part of this section discusses the results of the agent–based simulation on a square lattice network, detailing the experimental configurations for both global and local intervention schemes, and utilises figures from different perspectives to interpret the results.
@@ 4.1 Analytical results for well–mixed population
We present the core analytical findings regarding the maximization of the Expected Total Social Welfare ($SW$) with respect to the incentive $\theta$. The analysis is organized into two main scenarios: the zero-sum transfer ($a=1$) and the non-zero-sum transfer ($a\neq 1$). Furthermore, this section examines the profound influence of selection intensity ($\beta$) on the objective’s behaviour and proposes a numerical algorithm to approximate the optimal incentive for the case of inefficient transfer ($a<1$).
@@ 4.1.2 Case 2: Non–Zero Sum Transfer($a\neq 1$)
4.1.3 Influence of Selection Intensity ($\beta$)
  • Neutral Selection Limit ($\beta\rightarrow 0^{+}$)
  • Strong Selection Limit ($\beta\rightarrow+\infty$)
    $\lim_{\beta\rightarrow+\infty}\frac{f(x)}{g(x)}$ $=H_{N}+1$
@@ 4.1.4 Algorithm to Approximate the Optimal Incentive for Maximum Social Welfare
Below is the pseudocode of our simple local search algorithm (Algorithm 1) to find an approximation of the optimal incentive $\theta^{*}$ for maximizing social welfare.
@@
4.2 Numerical Simulation for well–mixed populations